Suppose M is between L and N. Use the segment addition postulate to solve for this variable. Then find the lengths of LM MN and LN

LM 7y+9

MN= 3y+4

LN = 143

thank you

Answer by nyc_function(2741)

In geometry, the segment addition postulate states that if B is between A and C, then AB + BC = AC.
We can apply this rule to answer your question.

L----------M----------N
7y + 9 + 3y + 4 = 143
10y + 13 = 143
10y = 143 - 13
10y = 130
y = 130/10
y = 13
=======================
LN = 7y + 9
LN = 7(13) + 9
LN = 100
=======================
MN = 3y + 4
MN = 3(13) + 4
MN = 39 + 4
MN = 43
======================
You were given LN to be 143. So, there's no need to find it.